A student was asked to prepare 1.0 L of a 2.0 M CuSO4 solution. The student placed 40.0 g of CuSO4 crystals in a volumetric flask and filled it with water to the 1000–mL mark. Was the resulting solution 2.0 M? Explain. 
Answer #1

Given weight of cuso4 W=40 g

Also volume of water V = 1000 ml

Molecular weight of cuso4 is =M.W= 160 g/ mol

Number of moles of cuso4 is n = W/M.W

=40g/160(g/mol)

= 0.25 mol

Molarity of a solution is = n/V=number of moles per litre

=0.25moles/1 litre

= 0.25 M

Hence molarity is 0.25 M